OpenStudy (anonymous):

the length of a rectagular mailing label is 4 centimeters less than twice the width. The perimeter is 52 centimeters. Find the dimensions of the label.

myininaya (myininaya):

L=2W-4 2W+2L=52

myininaya (myininaya):

2W+2(2W-4)=52 4W-8=52 4W=60 W=15 L=2(15)-4=30-4=26